# Discretion in the Loop: Human Expertise in Algorithm-Assisted College Advising

### v2

@@ 3.3.1 Impossibility Result on Identifying Expertise from Observational Data
The following proposition formalizes the nonidentifiability of expertly targeted action from observational data alone. The proof presents a counterexample of a pair of observationally equivalent SCMs that differ in whether Definition 3.1 is met.
Proposition 3.1 (Non-identifiability of expertly targeted action).
Expertly targeted action is impossible to detect from facts about the observational distribution over $\mathbf{X}_{t}$, $A_{t}$ and $Y_{t+1}$ alone.
@@
To determine whether expertise is present, we must establish both that $A_{t}$ affects $Y_{t+1}$ and that they are related by a common cause $U_{t}$ (Figure 3). In general, it is impossible to distinguish whether two variables are directly related, related by a common cause, or both using just observational data about those two variables [53]. We carefully construct two SCMs $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ (see complete proof in Appendix A.2.1). These SCMs induce identical observational distributions $\mathbb{P}(A_{t},Y_{t+1})$. $\mathcal{M}_{1}$ meets the definition of an expertly targeted action, while $\mathcal{M}_{2}$ does not. We achieve observational equivalence by tuning the exogenous noise variables in both models to produce the same distribution over $A_{t}$ and $Y_{t+1}$, even though their causal structures differ. ∎
Since it is impossible to identify expertly targeted action from observational data alone, qualitative data on how human decision-makers choose actions is crucial to understanding the deployment dynamics of algorithmic decision support systems. In §5.1, we describe how to use advisors’ qualitative logs of student interactions to supplement our understanding of whether and how they use non-algorithmic information to target their actions.
@@ 3.3.2 Necessary Statistical Condition for Expertise
We now consider a conditional dependence condition $A_{t}\not\perp Y_{t+1}|\textbf{X}_{t}$ that is a necessary condition for expertise to be present (under a further assumption). This condition is analogous to the conditional dependence condition used by Alur et al. [4]. [^14] Alur et al. [4] identify this conditional dependence condition as sufficient for expertise in the predictive case, since they define expertise as simply the presence of a causal influence of non-algorithmic information on predictions. However, this definition of expertise does not make sense in the interventional case; we should expect advisors’ actions to influence the outcomes positively to count as expertly taken. To establish the necessity of the conditional independence condition for expertly targeted action, we make the assumption that the SCM $\mathcal{A}$ is causally faithful.
Definition 3.2 (Causal faithfulness).
An SCM $\mathcal{M}$ is causally faithful when for all disjoint subsets of variables $\mathbf{X},\mathbf{Y},\mathbf{Z}\subseteq\mathbf{V}_{\mathcal{M}}$, $\mathbf{X}\perp\mathbf{Y}|\mathbf{Z}\implies\mathbf{X}\perp_{\mathcal{G}_{\mathcal{M}}}\mathbf{Y}\ |\ \mathbf{Z}$, where $\perp_{\mathcal{G}_{\mathcal{M}}}$ is the relation of d-separation relative to the model’s induced causal graph $\mathcal{G}_{\mathcal{M}}$.
The causal faithfulness condition is a standard non-degeneracy assumption, commonly made in causal inference. It is violated when causal effects perfectly cancel – in our case, if $A_{t}\leftarrow\mathbf{U}_{t}\to Y_{t+1}$ perfectly offsets $A_{t}\to Y_{t+1}$ – which is quite rare. [^15] For instance for linear models, it happens with zero probability when coefficients are drawn randomly [53, 68].
Proposition 3.2 (Necessary condition for expertise).
Assume causal faithfulness. If there is expertly targeted action, then $A_{t}\not\perp Y_{t+1}|\ \mathbf{X}_{t}$.
@@
For any of the three components of the expertly targeted action definition (Definition 3.1) to be met, it must be the case that $A_{t}\not\perp_{\mathcal{G}_{\mathcal{M}}}Y_{t+1}|\ \mathbf{X}_{t}$ (as Effective action requires $A_{t}\rightarrow Y_{t+1}$, Targeted action requires $U_{t}\rightarrow A_{t}$, and Heterogeneous action requires both). So then by causal faithfulness, $A_{t}\not\perp Y_{t+1}|\ \mathbf{X}_{t}$. ∎
Proposition 3.2 shows that, under causal faithfulness, a necessary condition for expertly targeted action is conditional dependence of the advisor’s action $A_{t}$ and the student’s outcome $Y_{t+1}|\mathbf{X}_{t}$. This proposition suggests a valuable method—by detecting conditional dependence, we detect a superset of actions exhibiting human expertise (putting power issues aside). We adapt the method from Alur et al. [4] to conduct this test in §4 and identify interventions meeting this necessary condition for expertise.
3.3.3 Sufficient Statistical Condition for Weakened Components of Expertise
Given the non-identifiability of expertly targeted actions (Proposition 3.1), the conditional dependence of actions $A_{t}$ and future outcomes $Y_{t+1}$ given $\mathbf{X}_{t}$ cannot be both a necessary and sufficient condition for expertly targeted actions. However, while insufficient to guarantee expertise, the statistical criterion can guarantee non-trivial components of it are met. To state the result, we define two concepts that weaken key components of the expertly targeted action definition, beginning with effective action:
@@ 3.3.3 Sufficient Statistical Condition for Weakened Components of Expertise
$\mathbb{P}(Y_{t+1}=1|do(A_{t}=1),\mathbf{X}_{t}=\mathbf{x},U_{t}=u)\neq\mathbb{P}(Y_{t+1}=1|do(A_{t}=0),\mathbf{X}_{t}=\mathbf{x},U_{t}=u)$
Impactful action guarantees that actions change the outcome for some individual, equivalent to the presence of edge $A_{t}\to Y_{t+1}$ (Figure 3.1). Effective actions, which improve the outcome, are a subset of impactful actions.
We can define a similar weakened criterion for targeting, guaranteeing that non-algorithmic context influences the advisor at some point in the decision-making process (Figure 3.2):
@@
$\mathbb{P}(A_{t}=1|do(U_{t}=1),\mathbf{X}_{t}=\mathbf{x})\neq\mathbb{P}(A_{t}=1|do(U_{t}=0),\mathbf{X}_{t}=\mathbf{x})$
Similar to impactful actions vs. effective action, non-algorithmic actions are a superset of targeted actions.
The following proposition demonstrates that meeting the statistical conditional dependence condition provides suggestive evidence for targeting, by showing that at least one of these necessary conditions of targeting are met. We defer its proof to Appendix A.2.2.
@@
If $A_{t}\not\perp Y_{t+1}|\ \mathbf{X}_{t}$, then at least one of impactful action (Definition 3.3) or non-algorithmic action (Definition 3.1.2) are met.
@@
Thus, if the statistical condition is met by an observational dataset, we can be certain that two necessary conditions for expertise are met: Within the entire population, there are individuals for whom the advisor’s actions are having an impact, are influenced by non-algorithmic information, or both. On the one hand, this provides another justification for using the condition to identify possible instances of expertly targeted action, without requiring a faithfulness assumption (Definition 3.2). On the other hand, the characterization in Proposition 3.3 brings to light several key limitations in interpreting the conditional independence condition; notably it does not establish whether the effect of the action on the outcome is heterogeneous with respect to $U_{t}$. [^16] As aforementioned, the condition does not guarantee expert targeting is present. Further, there is no guarantee that impactful action and non-algorithmic targeting are both met, or that they are met for the same individuals identified by $\mathbf{X}_{t}$ and $U_{t}$. The condition also does not confirm whether the action improves the chance of a good outcome, rather than merely changing those chances.
3.3.4 Implications for Expertise Auditing
@@
The theoretical framework suggests that testing for the conditional dependence of the actions $A_{t}$ and the outcomes $Y_{t+1}|\mathbf{X}_{t}$ allows us to identify interventions meeting a necessary condition of expertly targeted action, and a sufficient condition for weakened components of it. In the MAAPS dataset, this corresponds to testing for the conditional dependence between (1) advisors’ actions, as indicated by advisors in their meeting logs, and (2) students’ future outcomes, such as GPA or persistence, given (3) their algorithmically-available features. We show how to implement this test with our data in §4.
The theoretical framework also clarifies what is impossible to ascertain from quantitative observational data. In particular, since the condition is only sufficient for weakened components of expertly targeted action, the framework illuminates three important concerns to attend to in event of conditional dependence to determine whether expertly targeted action is genuinely present. These concerns guide our interpretation of our quantitative results (§4.2.1) as well as our qualitative analysis of advisor comments (§5).
(A) Non-Algorithmic Targeting.
Since the test only identifies a necessary condition for expertly targeted action, it is important to identify whether there is corroborating evidence that human decisionmakers draw on non-algorithmic information to guide their interventions, for instance using qualitative data about the decision-making process. This can substantiate the claim that $\mathbf{U}_{t}$, the non-algorithmic context, reflects non-algorithmic sources of information that advisors are incorporating into their interventions to improve students’ outcomes.
(B) Supporting At-Risk Students.
In many cases, it may be that advisors are targeting their supportive actions to students based on algorithmically-unobserved risk factors: pieces of information about the student that suggest they are at risk for poor outcomes but are not captured by features of risk prediction (for instance, an unpredictable family emergency). However, even if actions taken by advisors counterfactually help these students, it may appear as if they are correlated with negative outcomes because the actions are targeted to an especially at-risk population. Thus, identifying corroborating evidence for targeting based on risk factors can help disentangle two very different explanations of a test result suggesting such negative correlation: that actions are hurting students, or supporting at-risk students. [^17] It might be tempting to try to disentangle Impactful action (Definition 3.3) from Effective action (Definition 3.1.2) observationally by looking just for positive correlations between taking an action and an outcome. However, this strategy is untenable because it is always possible that especially at-risk students, even if helped through expertly targeted action, may perform worse than average, leading to negative correlation.
(C) Nuisance Confounding.
When pieces of non-algorithmic context $\mathbf{U}_{t}$ do not meet all the components of the expertly targeted condition, they may not be non-algorithmic information at all – instead, they may be nuisance confounders. For instance, if $A_{t}$ is an advisor making a schedule recommendation, and $Y_{t+1}$ is next-term enrollment, then the $\textbf{U}_{t}$ of a student requesting schedule advice is a nuisance confounder, since students who request such advice are disproportionately likely to enroll the next semester regardless of the advisor’s actions. It is important to identify any nuisance confounding that might drive conditional dependence, by considering the implications of variable choice in the statistical test and looking for qualitative evidence of nuisance confounding. [^18] The distinction between a nuisance confounder and non-algorithmic information requires judgment, and a variable might be viewed as playing both roles at once; for instance, it arguably takes expertise for advisors to recognize a student is asking for help when those cues might not always be so obvious as an explicit request. Given the subjective nature of the distinction, we do not formally represent the distinction between nuisance confounders and non-algorithmic information in the model.
4 Quantitative Findings: Empirical Investigation of Advisor Expertise
We test for the conditional independence condition given in Proposition 3.2, which we showed is a necessary condition for expert targeting and sufficient for weakened components of it. Based on the results, we find five actions with significant conditional dependence on students’ outcomes. This test suggests the possibility of expertly targeted interventions, although we remind the reader of the impossibility result on identifying human expertise in targeting from quantitative observational data alone (see §3.3.1). Further, we build on the discussion in §3.3.4 and consider whether, in the MAAPS setting, alternative explanations (other than expertly targeted actions) may explain the conditional dependence results (§4.2.3), motivating the qualitative analysis in §5.
@@ 4.1 ExpertTest*: A Conditional Independence Test for Interventions
To test for advisor expertise in the MAAPS experiment, we adapt the test proposed by Alur et al. [4], which audits for expertise in algorithm-assisted prediction, to our setting of algorithm-assisted interventions. We refer to this test as the ExpertTest*. [^19] Just as for the theoretical framework, in our setting of algorithm-aided intervention, we replace the predictions of the original ExpertTest’s human forecaster with the advisor’s actions at a time $t$, $A_{t}$ and the predictive target $Y$ with their outcomes at time $Y_{t+1}$. We include the asterisk because of two differences in our setting: First, when predictions/actions may influence the outcome, the test is insufficient to demonstrate expertise, and second, because unlike Alur et al. [4], we also test for negative association for reasons described below.
Given a dataset $\mathcal{D}$ of $n$ $(\mathbf{x}_{t}^{j},a_{t}^{i,j},y_{t+1}^{j})$ observations, instantiations of the random variables ($\mathbf{X}_{t},A_{t}^{i},Y_{t+1}$), we look for independence of actions $A_{t}$ and outcomes $Y_{t+1}$ given the available data to an algorithmic predictor, $\mathbf{X}_{t}$:
$H_{0}:A_{t}^{i}\perp Y_{t+1}|\mathbf{X}_{t}$
If the ExpertTest* fails, then advisor actions are predictive of the student’s future outcomes in ways that cannot be explained by administrative data alone—suggesting they could be targeted to the student using non-algorithmic information. Indeed, the ExpertTest* corresponds to the statistical condition considered in §3.3 that is a necessary condition for expertly targeted action (Definition 3.1), and a sufficient condition for weakened components of it. Moreover, the test identifies potential expertly targeted actions relative to any algorithm based on $\mathbf{X}_{t}$, enabling its application even when risk scores are not recorded (as is the case for the MAAPS study).
4.1.1 ExpertTest* Implementation
For completeness, we briefly describe the implementation of ExpertTest* following Alur et al. [4], and note one point of departure from the original method.
@@ 4.1.1 ExpertTest* Implementation
We start by constructing a sequence of $K+1$ datasets that would be exchangeable under the null hypothesis $H_{0}$ given above, the first of which $\mathcal{D}_{0}=\{(x_{t}^{1},a_{t}^{i,1},y_{t+1}^{i,1})\ldots(x_{t}^{n},a_{t}^{i,n},y_{t+1}^{i,n})\}$ is the true dataset, and the other $K$ of which are synthetic datasets $\mathcal{D}_{1}\ldots\mathcal{D}_{K}$. To construct those synthetic datasets, we identify $L$ pairs of individuals with (nearly) identical $\mathbf{X}_{t}$ values (“algorithmically-indistinguishable individuals”) in $\mathcal{D}_{0}$. [^20] When individuals in $\mathcal{D}_{0}$ cannot be paired exactly, as is the case for our dataset, the test can be relaxed to accommodate approximate pairing, introducing error that scales in match quality (measured by a distance function on $\mathbf{X}_{t}$). We refer readers to [4] for the justification of the relaxation as well as rigorous statistical justification of the test. Then, using those $L$ pairs, $K$ synthetic datasets are constructed from $\mathcal{D}_{0}$ by making swaps of the actions advisors take for paired individuals with probability $\frac{1}{2}$.
@@
The mean squared error of a dataset is defined as $MSE(\mathcal{D})=\sum_{j=1}^{n}(a_{t}^{i,j}-y_{t+1}^{j})^{2}$ with $n=|\mathcal{D}|$. Let $11[\alpha\gtrsim\beta]=$ 1 if $\alpha>\beta$ with ties broken randomly. Given exchangeability, the order statistics over the mean squared error $MSE(\mathcal{D})$ of these datasets,
@@
As discussed in §3.3.3, actions intended to assist at-risk students may be negatively associated with outcomes. This is a key point of departure from the pure prediction setting [4], where good predictions must be positively associated with the outcomes. To address this, for every action $A^{i+}_{t}$ (e.g., recommending a schedule), ExpertTest* is also applied to the transformed action variable $A_{t}^{i-}=1-A_{t}^{i+}$ (e.g., not recommending a schedule). This is equivalent to defining
@@
and running the same procedure described above using $MSE^{-}$, allowing us to detect cases where taking an action is associated with worse outcomes, consistent with effective targeting of hidden risks. [^21] Another reasonable choice would be to run a two-tailed test on the order statistic $\tau_{K}$, which we opt against primarily for reasons of interpretability. See Appendix A.3.2 for a comparison; the methods yield similar results.

### v3

@@ 4.3.2 Impossibility Result on Identifying Expertise from Observational Data
The following proposition formalizes the nonidentifiability of expertly targeted action from observational data alone. The proof, which is presented in full in Appendix A.2.1, presents a counterexample pair of observationally equivalent SCMs that differ in whether Definition 4.1 is met.
Proposition 4.1.
Suppose Assumptions 4.1 through 4.4 are met. Expertly targeted action is nonetheless impossible to detect from facts about the observational distribution over $\mathbf{X}_{t}$, $A_{t}$ and $Y_{t+1}$ alone.
@@
To determine whether expertise is present, we must establish first both that $A_{t}$ affects $Y_{t+1}$ and that they are related by a common cause $U_{t}$. But in general, it is impossible to distinguish whether two variables are directly related, related by a common cause, or both using just observational data about those two variables [69]. We carefully construct two SCMs $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ inducing identical observational distributions $\mathbb{P}(A_{t},Y_{t+1})$. [^12] We leave $\mathbf{X}_{t}$ out of both models without loss of generality. $\mathcal{M}_{1}$ satisfies the definition of expertly targeted action, while $\mathcal{M}_{2}$ does not. We achieve observational equivalence by tuning the exogenous noise variables in both models to produce the same distribution over $A_{t}$ and $Y_{t+1}$, even though their causal structures differ. ∎
Since it is impossible to identify expertly targeted action from observational data alone, in the absence of experimentation, qualitative data on how human decision makers choose actions is crucial to understanding the deployment dynamics of algorithmic decision aids.
@@ 4.3.3 Necessary Statistical Condition for Expertise
Proposition 4.2 shows that, under causal faithfulness (Assumption 4.4), a necessary condition for expertly targeted action is conditional dependence of the advisor’s action $A_{t}$ and the student’s outcome $Y_{t+1}|\mathbf{X}_{t}$.
Proposition 4.2.
Suppose the DAG induced by $\mathcal{A}$ meets the constraints in Figure 1(Assumption 4.1) and $\mathcal{A}$ is causally faithful (Assumption 4.4). Then $A_{t}\not\perp Y_{t+1}|\mathbf{X}_{t}$.
@@
For any of the three components of the expertly targeted action definition (Definition 4.1) to be met, it must be the case that $A_{t}\not\perp_{\mathcal{G}_{\mathcal{M}}}Y_{t+1}|\ \mathbf{X}_{t}$ (as effective action requires $A_{t}\rightarrow Y_{t+1}$, targeted action requires $U_{t}\rightarrow A_{t}$, and heterogeneous effects requires both). So then by causal faithfulness, $A_{t}\not\perp Y_{t+1}|\ \mathbf{X}_{t}$. ∎
This proposition suggests a valuable method—by detecting conditional dependence, we detect a superset of actions exhibiting human expertise (putting power issues aside).
4.3.4 Sufficient Statistical Condition for Weakened Related Criteria
While insufficient to guarantee expertise, the statistical condition can guarantee non-trivial related criteria are met. To state the result, we define criteria weakening the first two of the expertly targeted action definition, beginning with effective action:
@@ 4.3.4 Sufficient Statistical Condition for Weakened Related Criteria
$\mathbb{P}(Y_{t+1}=1|do(A_{t}=1),\mathbf{X}_{t}=\mathbf{x},U_{t}=u)$ $\neq$
$\mathbb{P}(Y_{t+1}=1|do(A_{t}=0),\mathbf{X}_{t}=\mathbf{x},U_{t}=u)$
Impactful action guarantees that actions change the outcome for some student. Effective actions, which improve the outcome, are a subset of impactful actions.
Targeted action can be similarly weakened to a criterion stating that non-algorithmic context changes the actions of the advisor for some algorithmically-indistinguishable students.
@@
$\mathbb{P}(A_{t}=1|do(U_{t}=1),\mathbf{X}_{t}=\mathbf{x})$ $\neq$
$\mathbb{P}(A_{t}=1|do(U_{t}=0),\mathbf{X}_{t}=\mathbf{x})$
Non-algorithmic actions are a superset of targeted actions.
The following proposition demonstrates that meeting the conditional dependence condition $A_{t}\not\perp Y_{t+1}|\ \mathbf{X}_{t}$ provides suggestive evidence for targeting, by showing that at least one of these criteria are met. We defer its proof to Appendix A.3.
@@
Suppose the DAG meets the constraints in Figure 1(Assumption 4.1), and both Assumptions 4.2 and 4.3 are satisfied. If $A_{t}\not\perp Y_{t+1}|\ \mathbf{X}_{t}$, then at least one of impactful action (Definition 4.2) or non-algorithmic action (Definition 4.3) is met.
@@
Thus, if an observational dataset meets the statistical condition, we can be certain that there are students for whom the advisor’s actions are having an impact, are influenced by non-algorithmic information, or both. On the one hand, this provides another justification for using the condition to identify possible instances of expertly targeted action, without requiring a strong faithfulness assumption (Assumption 4.4). On the other hand, it brings to light several key limitations in interpreting the conditional independence condition. In particular, it does not establish that the action’s effect is heterogeneous with respect to $U_{t}$, or whether the action improves the chance of a good outcome, rather than merely changing those chances. [^13] Further, it does not establish whether both impactful action and non-algorithmic targeting are met, or just one of the two, and that if both are met, they are met for the same individuals identified by $\mathbf{X}_{t}$ and $U_{t}$.
4.4 Implications for Expertise Auditing
@@
The theoretical framework suggests that testing for conditional dependence of advisor actions $A_{t}$ and outcomes $Y_{t+1}$ given algorithmically available features $\mathbf{X}_{t}$ allows us to identify interventions that meet a necessary condition of expertly targeted action (Proposition˜4.2) and a sufficient condition for weakened criteria (Proposition˜4.3). We implement this test on the MAAPS dataset in §5, discussing alternative explanations for a significant result in §5.4.
Though quantitative observational data cannot definitively identify expertise, qualitative data can be used to provide greater insight into advisor decision making and provide evidence for expertly targeted action above alternative explanations. In §6, we use advisor meeting notes to look for evidence of targeted action (Definition 4.1.2), advisors using non-algorithmic context to inform their choice to take an intervention.
5 Quantitative Findings: Empirical Evidence of Advisor Expertise
On paper, two students may appear indistinguishable, sharing standardized test scores, demographics, and financial aid statuses. Yet after meeting with each of them, an advisor might recommend different schedules—because the students differ in ways that are invisible from basic administrative data, such as their interests and career aspirations..
In this section, we describe results of a statistical test to identify situations like the example above in the MAAPS experiment’s treatment arm. When advisors target action to students using “non-algorithmic information” learned in advising meetings, these targeting decisions could not be captured by any possible algorithm based on administrative data alone. Correspondingly, we refer to actions that improve student outcomes because they were targeted based on non-algorithmic information as “expertly targeted actions” (formally defined in Definition 4.1), reflecting a uniquely human contribution to advising.
To search for evidence of expertly targeted action in the MAAPS dataset, we test for dependence of (1) advisor interventions recorded in meeting logs (denoted $A_{t}$) and (2) students’ future academic outcomes (denoted $Y_{t+1})$, like their GPA, given (3) algorithmically-available features about students (denoted $\mathbf{X}_{t}$):
$H_{0}:A_{t}\perp Y_{t+1}|\mathbf{X}_{t}$
Intuitively, the condition looks at whether algorithmically-available information about students exhausts association of advisor actions and student outcomes. If not, any remaining association may reflect the advisor drawing on non-algorithmic information relevant to the outcome to target their actions to otherwise indistinguishable students, like in the example vignette. [^14] For readers of the Theoretical Framework (§4), the test corresponds to the conditional dependence condition analysed in Propositions 4.2 and 4.3.
To implement the test, we adapt a method devised by Alur et al. [6], calling it the ExpertTest*. We describe the test and its application to our dataset in §5.1. Our tests find five interventions with significant results, reported in §5.2. However, the ExpertTest* has two key limitations. First, there are potential power limitations, as discussed in §5.3. Second, there are alternative explanations for conditional dependence than expertly targeted action, which we discuss in §5.4. Subsequent qualitative analysis in §6 provides further support for targeting based on non-algorithmic information.
@@ 5.1 ExpertTest*: A Conditional Independence Test for Interventions
To test for advisor expertise in the MAAPS experiment, we adapt the test proposed by Alur et al. [6], which audits for expertise in algorithm-assisted prediction, to our setting of algorithm-assisted interventions. We refer to this test as the ExpertTest*. [^15] Just as for the Theoretical Framework (§4), in our setting of algorithm-aided intervention, we replace the predictions of the original ExpertTest’s human forecaster with the advisor’s actions at a time $t$, $A_{t}$ and the predictive target $Y$ with their outcomes at time $Y_{t+1}$. We include the asterisk because of two differences in our setting: First, when predictions/actions may influence the outcome, the test is insufficient to demonstrate expertise, and second, because unlike Alur et al. [6], we also test for negative association for reasons described below.For completeness, we briefly describe the ExpertTest* following Alur et al. [6], and note one point of departure from the original method.
@@
We start by constructing a sequence of $K+1$ datasets that would be exchangeable under the null hypothesis $H_{0}$ given above, the first of which $\mathcal{D}_{0}=\{(x_{t}^{1},a_{t}^{i,1},y_{t+1}^{1})\ldots(x_{t}^{n},a_{t}^{i,n},y_{t+1}^{n})\}$ is the true dataset, and the other $K$ of which are synthetic datasets $\mathcal{D}_{1}\ldots\mathcal{D}_{K}$. To construct those synthetic datasets, we identify $L$ pairs of individuals with (nearly) identical $\mathbf{X}_{t}$ values (“algorithmically-indistinguishable individuals”) in $\mathcal{D}_{0}$. [^16] When individuals in $\mathcal{D}_{0}$ cannot be paired exactly, as is the case for our dataset, the test can be relaxed to accommodate approximate pairing, introducing error that scales in match quality (measured by a distance function on $\mathbf{X}_{t}$). We refer readers to [6] for the justification of the relaxation as well as rigorous statistical justification of the test. Then, using those $L$ pairs, $K$ synthetic datasets are constructed from $\mathcal{D}_{0}$ by making swaps of the actions advisors take for paired individuals with probability $\frac{1}{2}$. Then, using those $L$ pairs, $K$ synthetic datasets are constructed from $\mathcal{D}_{0}$ by making swaps of the actions advisors take for paired individuals with probability $\frac{1}{2}$.
@@
The mean squared error of a dataset is defined as $MSE(\mathcal{D})=\sum_{j=1}^{n}(a_{t}^{i,j}-y_{t+1}^{j})^{2}$ with $n=|\mathcal{D}|$. Let $1[\alpha\gtrsim\beta]=1$ if $\alpha>\beta$ with ties broken randomly. Given exchangeability, the order statistics over the mean squared error $MSE(\mathcal{D})$ of these datasets,
@@
The ExpertTest* only rejects the null whenever the mean squared error in the synthetic datasets is routinely higher than the the mean squared error of the original dataset. In Alur et al. [6], this one-sided test of predictions is sensible, as it ensures the test only detects instances where non-algorithmic information is used to improve the predictions of the human forecaster. But we are also interested in negative association, as actions intended to assist at-risk students may be negatively associated with outcomes due to selection bias for hidden risk factors (see §5.4). This is a key point of departure from the pure prediction setting [6], where good predictions must be positively associated with the outcomes. To address this, for every action $A^{i+}_{t}$ (e.g., recommending a schedule), ExpertTest* is also applied to the transformed action variable $A_{t}^{i-}=1-A_{t}^{i+}$ (e.g., not recommending a schedule). This is equivalent to defining
@@
and running the procedure described above using $MSE^{-}$, allowing us to detect cases where taking an action is associated with worse outcomes, consistent with effective targeting of hidden risks. [^17] Another reasonable choice would be to run a two-tailed test on the order statistic $\tau_{K}$, which we opt against primarily for reasons of interpretability. See Appendix A.5.2 for a comparison; the methods yield similar results.